In 1978 Amitsur and Saltman constructed generic abelian crossed product p-algebras containing no non-trivial p-power central elements. These algebras are examples of noncyclic p-algebras and the existence of such algebras answered a question of Albert. In this talk we discuss these terms and talk about why prime to p extensions of these algebras are of interest and what is known about them.